OpenStudy (anonymous):

@amistre64 For Function f(x)=7x-m/2-nx fine the values of find the values of m and n such that the vertical asymptote is at x=6 and the x-intercept is 5.

OpenStudy (anonymous):

I found the values but the thing is i did trial and error, and I wanted to know if there is a easier way to approach this?

OpenStudy (freckles):

\[f(x)=\frac{7x-m}{2-nx} \]

OpenStudy (freckles):

You can find x-intercepts by finding when the top is 0. 7x-m=0 You are given that x=5 is suppose to satisfy this equation so plug it in and solve for m.

OpenStudy (anonymous):

oh shoot wow, nvm lol i don't even know what happen to me there

OpenStudy (anonymous):

thanks

OpenStudy (freckles):

To find vertical asymptotes (or holes except this one won't have holes because the top factor won't cancel the bottom), you set bottom equal to 0 so we have 2-nx=0 and you know we have a va @ x=6 so replace x with 6 and solve for n.

OpenStudy (anonymous):

Got it m=35, and n=1/3

OpenStudy (anonymous):

Thanks again for getting me on track.

OpenStudy (freckles):

Np. Is that what you found by your trial and error?

OpenStudy (anonymous):

yes, and I did the easier way and got the same answer.